Prove that





lim 1/(x+5)^2 = ∞


x→-5

1
Expert's answer
2022-03-04T09:45:57-0500

Solution


To Prove


"\lim_{x\rightarrow -5}\frac{1}{(x+5)^{2}}=\infty"


For the function,


"f(x)=\frac{1}{(x+5)^{2}}"


The table of values is




From the table, we conclude that



"\lim_{x\rightarrow -5}\frac{1}{(x+5)^{2}}=\infty"


The same is shown by the following plot